# Non-relativistic transport from frame-indifferent kinetic theory

### v1

@@ IV.2 Currents and conservation laws
V Example: rotating gas in a gravitational field in equilibrium
In this section we construct the equilibrium distribution function of a rotating three-dimensional gas from the perspective of two different observers: a non-rotating one and a rotating one. This exercise serves two purposes. Firstly, it gives an opportunity to showcase the crucial role of the Milne boost in shifting between the rest frames of different observers. Secondly, it highlights the advantages of working with Milne-invariant data.
We assume that the gas is placed in a spherically symmetric potential. Furthermore, we also assume that the clock form is closed $d\tau=0$, meaning that there is a canonical coordinates system $x^{\mu}=(ct,x^{i})$, such that $\tau=dt$. These assumptions can be expressed in the cylindrical coordinate system $(t,z,\rho,\phi)$ by setting
$h_{\mu\nu}dx^{\mu}dx^{\nu}=d\rho^{2}+dz^{2}+\rho^{2}d\phi^{2},~{}~{}~{}~{}v^{\mu}\partial_{\mu}=\partial_{t},~{}~{}~{}~{}m_{\mu}dx^{\mu}=-\varphi(r)dt\,,$ (80)
where $r=\sqrt{z^{2}+\rho^{2}}$ and $\varphi(r)$ describes the external potential. A comparison with Eq. (5) reveals that this data corresponds to the rest frame of an observer with a non-zero geodesic acceleration $a^{\mu}\neq 0$, but a zero vorticity $\omega^{\mu\nu}=0$. Let us now consider a fluid with temperature $T$ and moving with a constant angular velocity $u^{\mu}=\delta^{\mu}_{0}+\Omega\delta^{\mu}_{3}$. It can be easily checked that the hydrostatic conditions in Eqs. (61) and (63) are satisfied. Using Eq. (40) we can then construct the Milne-invariant gauge field
$A_{\mu}dx^{\mu}=-\left(\varphi(r)+\frac{1}{2}\rho^{2}\Omega^{2}\right)dt+\rho^{2}\Omega\,d\phi\,.$ (81)
The hydrostatic conditions for the chemical potential given in Eq. (62) or, equivalently, in Eq. (183), become
$\partial_{z}\mu$ $=-\varphi^{\prime}(r)\hat{z}\,,$ (82)
$\partial_{\rho}\mu$ $=-\varphi^{\prime}(r)\hat{\rho}+\rho\Omega^{2}\,,$ (83)
$\partial_{\phi}\mu$ $=0\,,$ (84)
where $\hat{z}=z/r$ and $\hat{\rho}=\rho/r$. The solution is simply
$\mu=\mu_{0}-\varphi(r)+\frac{1}{2}\rho^{2}\Omega^{2}\,,$ (85)
where $\mu_{0}$ is an integration constant. This chemical potential $\mu$ in conjunction with the constant temperature $T$ and the velocity $u^{\mu}$ fully specifies the equilibrium distribution function, see Eqs. (52) and (54).
We will now want to describe the same system in the frame of reference rotating with the angular velocity $\Omega_{\mathrm{obs}}$, which can in general be different from the angular velocity of the gas $\Omega$. To this end, we first perform a change of variables
$d\phi=d\Phi+\Omega_{\mathrm{obs}}dt\,,$ (86)
where $\Omega_{\mathrm{obs}}$ will represents the relative angular velocity of the new frame with respect to the old frame. In the new coordinate system the form of $\tau_{\mu}$ and $h^{\alpha\beta}$ does not change, but the components of $h_{\alpha\beta}$ and $v^{\mu}$ do transform:
$h_{\mu\nu}dx^{\mu}dx^{\nu}=d\rho^{2}+dz^{2}+\rho^{2}(d\Phi+\Omega_{\mathrm{obs}}dt)^{2},~{}~{}~{}~{}v^{\mu}\partial_{\mu}=\partial_{t}-\Omega_{\mathrm{obs}}\partial_{\Phi},~{}~{}~{}~{}m_{\mu}dx^{\mu}=-\varphi(r)dt.$ (87)
We can return $h_{\alpha\beta}$ and $v^{\mu}$ to their more familiar form by performing a Milne boost given in Eqs. (1), (3) and (6) with $\psi_{\mu}dx^{\mu}=\rho^{2}\Omega_{\mathrm{obs}}d\Phi$. Then
$\begin{split}\left(h_{\psi}\right)_{\mu\nu}dx^{\mu}dx^{\nu}&=d\rho^{2}+dz^{2}+\rho^{2}d\Phi^{2},~{}~{}~{}~{}\left(v_{\psi}\right)^{\mu}\partial_{\mu}=\partial_{t},\\ \left(m_{\psi}\right)_{\mu}dx^{\mu}=&-\left(\varphi(r)-\frac{1}{2}\rho^{2}\Omega_{\mathrm{obs}}^{2}\right)dt+\rho^{2}\Omega_{\mathrm{obs}}d\Phi\,.\end{split}$ (88)
Notice that according to Eq. (5) $\left(v_{\psi}\right)^{\mu}$ exhibits a non-zero curl, $\omega^{\mu\nu}=4\rho\Omega_{\mathrm{obs}}\delta^{[\mu}_{2}\delta^{\nu]}_{3}$. It is now clear that both a coordinate change and a Milne boost are necessary to shift to the rest frame of a different observer, i.e., the frame in which $\left(v_{\psi}\right)^{\mu}=\delta^{\mu}_{0}$. The Milne boost makes it possible to preserve the standard tensorial form of the metric data $h_{\alpha\beta}$ and $v^{\mu}$ at the price of modifying the gauge field $m_{\mu}$. Such a modification gives rise to additional forces – in the present case they are the centrifugal force and the Coriolis force. This conforms with the standard treatment of non-inertial effects in classical mechanics.
In the rotating frame of reference, the velocity of the gas is $u^{\mu}=\delta^{\mu}_{0}+u_{\phi}\delta^{\mu}_{3}$, where $u_{\phi}=\Omega-\Omega_{\mathrm{obs}}$. After calculating $A_{\mu}$ and using Eq. (62) or Eq. (183), the hydrostatic conditions take the form
$\partial_{z}\mu$ $=-\varphi^{\prime}(r)\hat{z}\,,$ (89)
$\partial_{\rho}\mu$ $=-\varphi^{\prime}(r)\hat{\rho}+\rho\left(\Omega_{\mathrm{obs}}+u_{\phi}\right)^{2}\,,$ (90)
$\partial_{\phi}\mu$ $=0\,.$ (91)
Thus, we again obtain the chemical potential given by Eq. (85) with the replacement $\Omega=\Omega_{\mathrm{obs}}+u_{\phi}$.
At this point we can make three observations. Firstly, the formalism of the NC kinetic theory allows us to calculate the equilibrium distribution function in any frame of reference from first principles, without any ad-hoc assumptions as to the correct treatment of non-inertial effects.
Secondly, let us examine what effect a small angular velocity, $u_{\phi}\ll\Omega_{\mathrm{obs}}$, has on the equilibrium distribution of the gas from the perspective of a rotating observer. Denoting the gradient of the chemical potential in the radial direction for the non-rotating gas as $\partial_{\rho}\mu|_{u_{\phi}=0}=-\varphi^{\prime}(r)\hat{\rho}+\rho\Omega_{\mathrm{obs}}^{2}$, we have to the linear order in $u_{\phi}$ that $\partial_{\rho}\mu=\partial_{\rho}\mu|_{u_{\phi}=0}+2\rho\Omega_{\mathrm{obs}}u_{\phi}$. As we will show in the next section, the pressure of a gas is related to its density and temperature by $p=nT$ with $n\propto e^{\mu/T}$, meaning that the gradient of pressure necessary to sustain equilibrium changes as
$\partial_{\rho}p=\partial_{\rho}p|_{u_{\phi}=0}+2\Omega_{\mathrm{obs}}J_{\phi}\,,$ (92)
where $J_{\phi}\equiv n\rho u_{\phi}$ is the azimuthal current. This can be seen as an analogue of the Hall effect with the gradient of pressure playing the role of the electric field and the Coriolis force playing the role of the magnetic field.
Thirdly, we find that because of the centrifugal force, the chemical potential $\mu$ grows to infinity with the distance to the axis of rotation $\rho$. This means that a rotating gas in an infinite space cannot, strictly speaking, exist in equilibrium. On the other hand, the equilibrium distribution can still be used when its regime of validity is restricted to a finite region of space. As an aside, we note that when trying to find the consequences of this observation for rotating planets one has to bear in mind that the atmosphere of a planet is an out-of-equilibrium system even in the absence of rotations because of the processes of atmospheric escape [70], so the conclusions of this section do not apply directly.
@@ VI NC hydrodynamics from kinetic theory
One of the most important roles of kinetic theory is that it provides a microscopic foundation for hydrodynamics of weakly coupled gases. Therefore, in order to lay the groundwork for future developments and uses of our formalism, in this section we derive equations of Newton-Cartan hydrodynamics and fix the first-order transport coefficients in two- and three-dimensional gases. The results are compared to analogous non-covariant results scattered in the literature.
@@ VI.1 Zeroth-order hydrodynamics
From this section onwards, we specify to the torsionless case $H_{\mu\nu}=0$ for simplicity. Importantly, this also removes the distinction between the different covariant derivatives, so that ¯$\bar{\nabla}_{\mu}=\nabla_{\mu}$.
@@
$\mathcal{T}^{\mu\nu}_{0}$ $=mnu^{\mu}u^{\nu}+nTh^{\mu\nu}\,,$ (94)
$\mathcal{E}^{\mu}_{0}$ $=\frac{dnT}{2}u^{\mu}\,.$ (95)
@@
Let us now look at near-equilibrium states with a distribution function that can be expressed as $f(x,p)=f_{0}(x,p)+\delta f(x,p)$. Analogically, we split the moments of the distribution function $f(x,p)$ into $\mathcal{I}^{\alpha_{1}\alpha_{2}\cdots\alpha_{k}}=\mathcal{I}^{\alpha_{1}\alpha_{2}\cdots\alpha_{k}}_{0}+\delta\mathcal{I}^{\alpha_{1}\alpha_{2}\cdots\alpha_{k}}$, with an explicit formula for $\mathcal{I}^{\alpha\beta\cdots\gamma}_{0}$ discussed below Eq. (95). A certain convention has to be chosen to define the out-of-equilibrium $n$, $u^{\mu}$ and $T$ that specify the ideal part of the distribution; we choose to define the out-of-equilibrium density, velocity and temperature using the zeroth, first, and the trace of the second moment of $f(x,p)$:
$n\equiv\mathcal{I},~{}~{}~{}~{}nu^{\mu}\equiv\mathcal{I}^{\mu},~{}~{}~{}~{}dnT\equiv mg_{\mu\nu}\mathcal{I}^{\mu\nu}\,.$ (96)
This choice corresponds to the Eckart frame [71, 57]. Note that all the quantities $n$, $T$, $u^{\mu}$ are Milne-invariant. In this work, we choose to express the collision term within the relaxation time approximation as
$\mathcal{C}[f]=-\frac{\delta f}{\tau_{\mathrm{coll}}}\,.$ (97)
With this choice we have
$\mathcal{K}^{\alpha_{1}\cdots\alpha_{k}}=-\frac{1}{\tau_{\mathrm{coll}}}\delta\mathcal{I}^{\alpha_{1}\alpha_{2}\cdots\alpha_{k}}\,,$ (98)
and therefore the identities $\mathcal{K}=\mathcal{K}^{\mu}=g_{\mu\nu}\mathcal{K}^{\mu\nu}=0$, see Eq. (76), are automatically satisfied. For the number current, momentum current and energy current we write
$\mathcal{J}^{\mu}=$ $nu^{\mu}\,,$ (99)
$\mathcal{T}^{\mu\nu}=$ $nTh^{\mu\nu}+mnu^{\mu}u^{\nu}+\tau^{\mu\nu}\,,$ (100)
$\mathcal{E}^{\mu}=$ $\frac{dnT}{2}u^{\mu}+q^{\mu}\,,$ (101)
where $\tau^{\mu\nu}\equiv m\delta\mathcal{I}^{\mu\nu}$ is traceless, symmetric, Milne-invariant, and spatial: $\tau_{\mu}\tau^{\mu\nu}=0$, while $q^{\mu}\equiv\frac{m}{2}g_{\nu\lambda}\delta\mathcal{I}^{\mu\nu\lambda}$ is Milne-invariant and spatial: $\tau_{\mu}q^{\mu}=0$. With these definitions, the conservation equations presented in Eqs. (77), (78) and (79) can be rewritten as equations of motion for $n$, $u^{\nu}$ and $T$:
$u^{\mu}\partial_{\mu}n+n\nabla_{\mu}u^{\mu}=0\,,$ (102a)
$mnu^{\mu}\nabla_{\mu}u^{\nu}+h^{\mu\nu}\partial_{\mu}p+\nabla_{\mu}\tau^{\mu\nu}=0\,,$ (102b)
$\frac{d}{2}nu^{\mu}\partial_{\mu}T+nT\nabla_{\mu}u^{\mu}+\tau^{\mu\nu}g_{\rho\mu}\nabla_{\nu}u^{\rho}+\nabla_{\mu}q^{\mu}=0\,,$ (102c)
where the pressure $p$ is
$p=nT\,.$ (103)
Furthermore, the general equation of motion for the moments given in Eq. (73) can be rewritten as
@@
The term $X_{L}[f_{0}]$ was already calculated in Eq. (59), where it was found to be proportional to $\delta_{\chi}h^{\mu\nu}$, $\delta_{\chi}A_{\mu}$ and $\delta_{\chi}\tau_{\mu}$. This explicitly proves that $\mathcal{S}^{\alpha\beta\ldots\gamma}$ vanishes when evaluated on the Killing data.
Equations (102) contain two unknown tensors $\tau^{\mu\nu}$ and $q^{\mu}$ which have to be determined. In the hydrodynamic approach, this is done by expanding $\tau^{\mu\nu}$ and $q^{\mu}$ in spatial gradients of the hydrodynamic variables $T$, $\mu$ and $u^{\mu}$. A gradient expansion that generalizes the Chapman-Enskog method to a NC gas can be performed if the timescales of out-of-equilibrium perturbations are much larger than the timescales at which non-conserved quantites relax, so that local equilibration can be assumed. Due to the rather technical character of these calculations, we relegate the details of the gradient expansion procedure, as well as a more thorough discussion of the assumptions involved, to Appendix A.

### v2

@@ IV.2 Detailed balance and thermal equilibrium
V Rotating gases: rest frame and equilibrium
Figure 1: Shifting to the reference frame co-rotating with the fluid, as described in Sec. V. The worldlines of the fluid elements (red dashed lines) at a fixed $\rho$ are plotted for different coordinate systems: $(t,\phi)$ and $(t,\phi^{\prime})$, and different choices of the reference velocity: $v^{\mu}$ and $v^{\prime\mu}$. The tilt of the $t$ axis is given by the angle between $\partial_{t}$ and $\partial_{\phi}$ (left column) or $\partial_{\phi^{\prime}}$ (right column) calculated using the metric $h_{\mu\nu}$ (top row) or $h^{\prime}_{\mu\nu}$ (bottom row).
The primary goal of this section is to highlight the crucial role played by the Milne boost in shifting between the rest frames of different observers. For simplicity, we will consider a two-dimensional NC spacetime [^7] The discussion straightforwardly generalizes to the three-dimensional case. of the form (18) deformed by a static rotationally invariant scalar potential. In addition, we will assume a gas rotating rigidly on such a geometry with angular velocity $\Omega$. In polar coordinates $(t,\rho,\phi)$, the fluid velocity reads
$u^{\mu}\partial_{\mu}=\partial_{t}+\Omega\partial_{\phi}\,.$ (86)
The Milne invariant fields for such spacetime are $\tau=dt$ and $h^{\mu\nu}\partial_{\mu}\otimes\partial_{\nu}=\partial_{\rho}\otimes\partial_{\rho}+\rho^{-2}\partial_{\phi}\otimes\partial_{\phi}$. In the first frame to be considered, the Milne dependent fields read
$h_{\mu\nu}dx^{\mu}dx^{\nu}=d\rho^{2}+\rho^{2}d\phi^{2},~~~~v^{\mu}\partial_{\mu}=\partial_{t},~~~~m_{\mu}dx^{\mu}=-\varphi(\rho)dt\,.$ (87)
Notice that relations (6) imply that the vorticity of such family of observers vanishes, whereas their geodesic accelaration is $a_{v}^{\mu}=\varphi^{\prime}\delta^{\mu}_{1}$.
On the other hand, we notice that by using the Milne parameter $\psi_{\mu}dx^{\mu}=\rho^{2}\Omega d\phi$, we can change the frame such that the new reference velocity is $v^{\prime\mu}=u^{\mu}=v^{\mu}+h^{\mu\nu}\psi_{\nu}$. By construction, on such a frame the geodesic accelaration $a_{v^{\prime}}^{\mu}$ and vorticity $\omega_{v^{\prime}}^{\mu\nu}$ will coincide with the ones of the fluid. In the co-moving frame, the primed Newton-Cartan data reads
$h^{\prime}_{\mu\nu}dx^{\mu}dx^{\nu}=d\rho^{2}+\rho^{2}(d\phi-\Omega dt)^{2},~~~~v^{\prime\mu}\partial_{\mu}=u^{\mu}\partial_{\mu},~~~~m^{\prime}_{\mu}dx^{\mu}=-\left(\varphi(\rho)-\frac{1}{2}\rho^{2}\Omega^{2}\right)dt+\rho^{2}\Omega\left(d\phi-\Omega dt\right)\,.$ (88)
However, the worldlines of the fluid elements are not given by paths with a constant $\phi$ since $u^{\mu}\partial_{\mu}\phi\neq 0$, see Fig. 1. To define the rest frame of the fluid, we require that in addition to adjusting the reference velocity such that $v^{\mu}=u^{\mu}$, the coordinate system has to be also adjusted such that $u^{\mu}\partial_{\mu}x^{\nu}=\delta^{\nu}_{0}$. In our problem, after performing the change of coordinates $t^{\prime}=t$, $\rho^{\prime}=\rho$, and $\phi^{\prime}=\phi-\Omega t$, the basis $\partial_{\rho^{\prime}},\partial_{\phi^{\prime}}$ gets locked to the motion of the gas, and the NC fields become
$h^{\prime}_{\mu\nu}dx^{\prime\mu}dx^{\prime\nu}=d\rho^{\prime 2}+\rho^{\prime 2}d\phi^{\prime 2},~~~~v^{\prime\mu}\partial^{\prime}_{\mu}=\partial_{t^{\prime}},~~~~m^{\prime}_{\mu}dx^{\prime\mu}=-\left(\varphi(\rho)-\frac{1}{2}\rho^{2}\Omega^{2}\right)dt^{\prime}+\rho^{2}\Omega d\phi^{\prime}\,.$ (89)
After calculating the field strength $F^{\prime}_{\mu\nu}=\partial_{\mu}m^{\prime}_{\nu}-\partial_{\nu}m^{\prime}_{\mu}$, the geodesic accelaration and vorticity can be found as
$a_{v^{\prime}}^{\mu}$ $=(\varphi^{\prime}-\rho^{\prime}\Omega^{2})\delta_{1}^{\mu}\,$, (90)
$\omega_{v^{\prime}}^{\mu\nu}$ $=2\Omega\,u^{\alpha}\epsilon_{\alpha}\,^{\mu\nu}\,.$ (91)
Remarkably, this geometric formulation naturally encodes the similarity of the non-inertial effects (the so-called centrifugal and Coriolis forces) to the Lorentz force of electrodynamics. Notice also that if the scalar potential is harmonic such that $\varphi=1/2\Omega\rho^{2}$, the centrifugal force will cancel with the central force and the accelaration vanishes.
This example illustrates how the presence of the gas naturally allows us to extend the original NC structure into the thermal NC spacetime introduced in Sec. II.2 using the fluid velocity. In fact, the Milne invariant fields $g_{\mu\nu},A_{\mu}$ correspond to $h^{\prime}_{\mu\nu},m^{\prime}_{\mu}$, respectively, in our construction. Therefore, in the thermal NC language, the fluid rest frame is given by the geometric data $(\tau_{\mu},u^{\mu},h^{\mu\nu},g_{\mu\nu},A_{\mu})$ expressed in coordinates such that $u^{\mu}=\delta^{\mu}_{0}$.
In addition, since the spacetime considered here possesses Killing vectors $\partial_{t},\partial_{\phi}$, and the fluid velocity is a linear combination of them, $u^{\mu}$ is a Killing vector. A straightforward calculation reveals that the hydrostatic conditions in Eqs. (83)-(85) are satisfied if the chemical potential and temperature are
$T$ $=\text{const}\,$, (92)
$\mu$ $=-\varphi(\rho)+\frac{1}{2}\rho^{2}\Omega^{2}+\text{const}\,.$ (93)
This relation implies that in flat space, finite particle density at infinity requires the scalar potential to diverge as $1/2\Omega\rho^{2}$, with the fine-tuned case $\varphi=1/2\Omega\rho^{2}$ corresponding to an equilibrium state with uniform density. Contrary to relativistic gases where causality forbids equilibrium states with rigid rotation, non-relativistic gases have no such formal obstruction, although such equilibrium can be considered unphysical due to the divergence of the particles’ density at the spatial infinity.
@@ VI NC hydrodynamics from kinetic theory
One of the most important contributions of kinetic theory is that it provides a microscopic foundation for hydrodynamics of weakly coupled gases. Therefore, in order to lay the groundwork for future developments and uses of our formalism, in this section we derive equations of Newton-Cartan hydrodynamics and fix the first-order transport coefficients in two- and three-dimensional gases. The results are compared to analogous non-covariant results scattered in the literature.
@@ VI.1 Zeroth-order hydrodynamics
From this section onwards, we specify to the torsionless case $H_{\mu\nu}=0$ for simplicity. Importantly, this also removes the distinction between the different covariant derivatives, so that $\bar{\nabla}_{\mu}=\nabla_{\mu}$.
@@
[…]
where the mass density, internal energy and pressure are $\rho=mn$, $\epsilon=\frac{dnT}{2}$ and $\mathfrak{p}=nT$ respectively. On the other hand, the ideal constitutive relations for an observer with velocity $v^{\mu}$ are
$\mathcal{P}_{\mu}$ $=\rho\,u_{\mu}\,,\qquad T_{\mu\nu}=\mathfrak{p}\,h_{\mu\nu}\,$, (96)
$\tilde{\mathcal{E}}^{\mu}$ $=\left(\epsilon+\frac{\rho}{2}u^{2}\right)u^{\mu}+\mathfrak{p}\,P^{\mu}_{\nu}\,u^{\nu}\,.$ (97)
@@
As previously argued, $f_{0}$ solves the Boltzmann equation only in the hydrostatic regime. Therefore, Eqs. (95) are exact only at equilibrium. However, since hydrodynamics is based on the idea of local equilibrium, it is natural to expand the out-equilibrium distribution function around the hydrodynamic distribution $f_{0}(x,p)$ evaluated at non-hydrostatic velocity, temperature and chemical potential defined as in (77) and (79), so that $f(x,p)=f_{0}(x,p)+\delta f(x,p)$. Consequently, we split the moments of the distribution function $f(x,p)$ into $\mathcal{I}^{\alpha_{1}\alpha_{2}\cdots\alpha_{k}}=\mathcal{I}^{\alpha_{1}\alpha_{2}\cdots\alpha_{k}}_{0}+\delta\mathcal{I}^{\alpha_{1}\alpha_{2}\cdots\alpha_{k}}$. In addition, we restrict the analysis to systems where the relaxation time approximation is applicable
$\mathcal{C}[f]=-\frac{\delta f}{\tau_{\mathrm{coll}}}\,$, (98)
with $\tau_{\mathrm{coll}}$ a phenomenological parameter that by construction is a scalar. With this choice we have
$\mathcal{K}^{\alpha_{1}\cdots\alpha_{k}}=-\frac{1}{\tau_{\mathrm{coll}}}\delta\mathcal{I}^{\alpha_{1}\alpha_{2}\cdots\alpha_{k}}\,.$ (99)
However, (covariant) conservation of particle number, momentum and energy requires $\delta\mathcal{I}=\delta\mathcal{I}^{\mu}=g_{\mu\nu}\delta\mathcal{I}^{\mu\nu}=0$, conditions that are automatically satisfy if we fix the Eckart frame [72, 58]. The ambiguity in the definition of the hydrodynamic variables ($\mu,T,u^{\mu}$) is fixed by requiring that the equations of state and the particle current do not receive higher derivative corrections
$n=c\mathcal{N}\left(2\pi mT\right)^{d/2}e^{\mu/T}\,,\qquad\epsilon=\frac{d}{2}nT\,,\qquad\mathcal{J}^{\mu}=nu^{\mu}\,.$ (100)
Notice that in the discussion of hydrodynamic frames, the emphasis is usually placed on fixing the velocity. However, we stress that it is essential to fix all thermodynamic functions that are not uniquely defined out of equilibrium.
The out-of-equilibrium currents then can be expressed as
$\mathcal{T}^{\mu\nu}=$ $\rho u^{\mu}u^{\nu}+\mathfrak{p}\,h^{\mu\nu}+\tau^{\mu\nu}\,$, (101)
$\mathcal{E}^{\mu}=$ $\epsilon\,u^{\mu}+q^{\mu}\,$, (102)
with the higher order corrections to the currents a transverse tensor a vector respectively, defined as $\tau^{\mu\nu}\equiv m\delta\mathcal{I}^{\mu\nu}$ $q^{\mu}\equiv\frac{m}{2}g_{\nu\lambda}\delta\mathcal{I}^{\mu\nu\lambda}$. In addition, the correction to the stress-momentum tensor is traceless $\tau^{\mu\nu}g_{\mu\nu}=0$. With these definitions, the conservation equations (69), (70), and (71), can be rewritten as equations of motion for $\mu$, $u^{\nu}$, and $T$:
$u^{\mu}\partial_{\mu}n+n\nabla_{\mu}u^{\mu}=0\,$, (103)
$u^{\mu}\nabla_{\mu}u^{\nu}+h^{\mu\nu}\partial_{\mu}\mathfrak{p}+\nabla_{\mu}\tau^{\mu\nu}=0\,$, (104)
Furthermore, the general equation of motion for the moments is given in Eq. (65) can be rewritten as
@@
The term $X_{L}[f_{0}]$ was already calculated in Eq. (81), where it was found to be proportional to $\delta_{\chi}h^{\mu\nu}$, $\delta_{\chi}A_{\mu}$ and $\delta_{\chi}\tau_{\mu}$. This explicitly proves that $\mathcal{S}^{\alpha\beta\ldots\gamma}$ vanishes when in equilibrium.
Equations (103)-(105) contain two unknown tensors, $\tau^{\mu\nu}$ and $q^{\mu}$, which have to be determined in order for the system of equations  (103)-(105) to be solvable beyond the ideal order. In the hydrodynamic approach, this is done by expanding $\tau^{\mu\nu}$ and $q^{\mu}$ in spatial gradients of the hydrodynamic variables $T$, $\mu$ and $u^{\mu}$. A gradient expansion that generalizes the Chapman-Enskog method to a NC gas can be performed if the timescales of out-of-equilibrium perturbations are much larger than the timescales at which non-conserved quantities relax, so that local equilibration can be assumed. Due to the rather technical character of these calculations, we relegate the details of the gradient expansion procedure, as well as a more thorough discussion of the assumptions involved, to Appendix A.